A long straight vertical wire carries an upward of Need help on how to solve this problem (3.4#1)
Answer #1

The magnetic field at AB due to current 10A in he straight wire, B(AB)=\mu0 * I /2\Pi r =4π × 107 *10/2π*0.01 =2× 10−4

Force on AB ,F=ILBsin\Theta =5*0.5*2× 10−4 *sin 90=5× 10−4 N

The direction is into the paper

similarly the field at CD,   B(CD)=\mu0 * I /2\Pi r =4π × 107 *10/2π*0.03 =0.667× 10−4

Force on AB ,F=ILBsin\Theta =5*0.5* 0.667× 10−4 *sin 90=1.667× 10−4 N

If we take the torque about CD 5× 10−4 N*(0.3-0.1)=10−4 N.m and tends to rotate the loop about CD into the page